These were accepted.
Stephen will continue to send out the report with the names of the machines requiring rebooting. Ongoing - but should make the 220 kernels a priority.
This has been amended to ask all Units to identify bids for old kit.
This has been amended. We agreed to defer removing machines until Ian has checked that AT server room.
Craig has emailed Angus Rae
Providing there is sufficient time, this will be scheduled after Alastair's documentation talk on 30th January.
All students that were asked had received notification although some thought that there were some students who hadn't.
See discussion points below.
Blog articles suggested:
Report from Computing Executive Group
The new documentation site should go live on the Monday following documentation week i.e. Monday 18th February. There will be a demo at 10am on Wednesday 30th January in 1.16.
Reports from units
Further work is scheduled at Kings Buildings for Saturday 26th January. There are a couple of staffmail outages associated with this work planned for Thursday and Monday. Subsequent to the meeting, George emailed sys-announce to advise users.
The annual routine maintenance on the UPSs is being carried out on Thursday 31st January.
The Infrastructure Unit have completed their SL6 upgrades.
Stephen advised that with a small server header, the default disk size of 15GB is likely to be more than sufficient in most cases. In other cases, Units need to make an informed decision on basis of requirements. Ian asked whether it is possible to make dynamic changes to disk partition sizes This isn't feasible.
IS have now produced an Office365 Guide for Apple Mail which includes IMAP instructions which is linked to from the Office365 FAQ.
The first student CO clinic is scheduled for Wednesday 30th January in the level 4 open area, AT, from 2:30 - 3:30. Need names of volunteers from each Unit (Graham has offered to be RAT representative).
Group 1 Craig, Ian, Iain, Jennifer, Richard will look into running one of the touch panels off a raspberry pi.
Group 2 Neil, Alastair, Toby, Ross will be meeting again soon to decide.
Group 3 Roger, Chris, George, Graham, Lindsey will be meeting again soon to decide.
Group 4 Gordon, Stephen, Tim, Alison, Carol will be looking at content on the touchscreens e.g. displaying nagios information.
It was agreed that it would not be a problem if groups chose the same topic but it doesn't look as though this will be the case this time round.
Topics for discussion
The following services were suggested: rfe, nsu, svn
Iain suggested that we fill in the telephone survey. See email from Chris Adie, sent 21st January.
The next meeting will be on Wednesday 13th February (documentation week!) at 10am in IF-4.31.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh